THE (1)^{1}Π_{g} STATE OF ^{39}K_{2} REVISITED

I. RUSSIER, M. AUBERT-FRECON, \underline{A. J. ROSS}, F. MARTIN, A. YIANNOPOULOU^{a}, and P. CROZET, Laboratoire de Spectrométrie Ionique et Moléculaire (CNRS UMR5579), and Université Lyon I, Campus la Doua, Bâtiment 205, 69622 Villeurbanne Cedex, France.

Vibrational levels up to v=107 of the (1)^{1}Π_{g} state of^{39}K_{2}, which dissociates to K(4s)^{2}S+K(4p)^{2}P_{3/2}, have been observed in high resolution Fourier transform records of C^{1}Π_{u} → (1)^{1}Π_{g} fluorescence of potassium molecules in a heatpipe source. From a revised analysis of this system, we have constructed a rotationless RKR potential curve which extends to R=40.3 Å covering 99.2% of the (1)^{1}Π_{g} state. Vibrational and rotational constants (at J=0) have been calculated from this curve. The parameters describing the coulombic interactions between K(4s) + K(4p) atoms are given in a Hund’s case (c) basis. From this work, the dissociation energy D_{e} for the (1)^{1}Π_{g} state is found to be 1290.30±0.09 cm^{−1} (leading to a new value of T_{e} = 16203.25 ± 0.09 cm^{−1} for this state) and the long-range parameters are M^{2} = 25.30 ± 0.02 e^{2}a_{o}^{2}, R_{01} = 7261 ± 223 e^{2}a_{o}^{5}. These values are in good agreement with values obtained from the pure long range 0_{g}^{−} state analyzed by H. Wang, P. L. Gould and W. C. Stwalley (J. Chem. Phys. 106, 7899 (1997)).

aCurrent address: Laboratoire Aimé Cotton (CNRS UPR3321), Campus d’Orsay, Bâtiment 505, 91405 Orsay, France